# SwiLTra-Bench: The Swiss Legal Translation Benchmark

### v1

@@ Abstract
In Switzerland legal translation is uniquely important due to the country’s four official languages and requirements for multilingual legal documentation. However, this process traditionally relies on professionals who must be both legal experts and skilled translators—creating bottlenecks and impacting effective access to justice. To address this challenge, we introduce SwiLTra-Bench, a comprehensive multilingual benchmark of over 180K aligned Swiss legal translation pairs comprising laws, headnotes, and press releases across all Swiss languages along with English, designed to evaluate LLM-based translation systems. Our systematic evaluation reveals that frontier models achieve superior translation performance across all document types, while specialized translation systems excel specifically in laws but under-perform in headnotes. Through rigorous testing and human expert validation, we demonstrate that while fine-tuning open SLMs significantly improves their translation quality, they still lag behind the best zero-shot prompted frontier models such as Claude-3.5-Sonnet. Additionally, we present SwiLTra-Judge, a specialized LLM evaluation system that aligns best with human expert assessments.111The datasets and code are available at https://huggingface.co/collections/joelniklaus/swiltra-bench-67c569a2ada47e4549733deb
@@
Joel NiklausH Jakob MeraneE,L,M Luka Nenadic E Sina Ahmadi Z Yingqiang Gao Z Cyrill A. H. Chevalley B Claude Humbel Z Christophe Gösken E Lorenzo Tanzi G Thomas Lüthi S Stefan Palombo H Spencer Poff H Boling Yang H Nan Wu H Matthew Guillod H Robin Mamié C Daniel Brunner C Julio Pereyra H Niko Grupen H HHarvey EETH Zurich CSwiss Federal Supreme Court ZUniversity of Zurich BUniversity of Basel GUniversity of Geneva LUniversity of Lausanne SCanton of Solothurn MMax Planck Institute for Research on Collective Goods
@@ 2.2 Dataset Splits
We first segment each dataset by a unique identifier (entire laws, entire headnotes, and entire press releases) to ensure that no single law, headnote, or press release is split across training, validation, and test. For laws, we prioritize examples for the validation and test splits that (1) have more language versions (to guarantee good multilingual coverage), (2) have an official abbreviation (since abbreviations are only set for those laws that are presumed to be cited frequently222https://www.bk.admin.ch/apps/gtr/de/index.html, which we consider a good proxy for practical importance, (3) have shorter text lengths to make evaluation faster and cheaper, and (4) have newer applicability dates so that more recent and multilingual laws are prioritized for validation and testing, resulting in a more realistic evaluation setting. For headnotes, we similarly prioritize those with more recent publication years for validation and test. Finally, for press releases, we focus on maximizing multilingual coverage by ensuring all validation and test examples are available in all present languages (German, French and Italian). The training sets contain all examples not held out for validation or testing.
@@ 2.3 Data Statistics
Table 1 presents the overall statistics of the three datasets included in SwiLTRa-Bench. We visualize the training set text lengths for the shortest levels used for training and evaluation in Figure B.1. For completeness, we show histograms for all levels in Figure B.2 and Figure B.3. To calculate these statistics, we used an NLTK333https://www.nltk.org word tokenizer, splitting sentences based on whitespace and punctuation.
@@ 3.1 Evaluation
To paint a representative picture of translation capabilities, we evaluate models across five main classes: 1) translation models, i.e., models specifically trained for translation tasks, 2) frontier models, i.e., large foundation models pre-trained on web-scale data and post-trained on diverse tasks, 3) reasoning models, i.e., models using significant resources at test time to improve output quality, 4) open models, i.e., typically small language models (SLMs) with publicly available weights, and 5) fine-tuned models, i.e., models specifically fine-tuned on SwiLTRa-Bench. We conducted our evaluation using the lighteval framework due to its ease of use and good support for custom metrics.444https://github.com/huggingface/lighteval
@@ 3.2 Fine-tuning
To provide an overview of the current open SLM landscape, we fine-tuned Gemma-2 2B and 9B Team et al. (2024), Llama 1B, 3B and 8B Grattafiori et al. (2024), Phi-3.5 mini and Phi-3 medium Abdin et al. (2024), and Qwen2.5 0.5B, 1.5B, 3B, 7B, 14B and 32B Team (2024) models on our dataset. We fine-tuned using Hugging Face transformers555https://github.com/huggingface/transformers and unsloth666https://github.com/unslothai/unsloth using 4-bit quantization and 8bit AdamW Loshchilov and Hutter (2019); Dettmers et al. (2022) on a single 80GB NVIDIA H100 GPU. We used rank stabilized LoRA Hu et al. (2021); Kalajdzievski (2023) with rank 16 and alpha 16. We trained with the model’s native chat template on sequence length 512, covering more than 99% of the training dataset and truncating the rest. We used packing, weight decay 0.01, batch size 128 and early stopping with patience 3. In most cases, the lowest evaluation loss is reached after exactly 1 epoch. We used a linear learning rate schedule with 1000 warmup steps and learning rate . We manually tuned the learning rate ( - ), weight decay (0.01, 0.1), label smoothing (factor 0, 0.01, 0.1) and LoRA rank (16, 128). We used the train and validation sets of the Law and Headnote translations on the lowest (shortest) levels, i.e., the paragraph and text levels. For all fine-tuned models, we used the instruction-tuned variant since they have shown to better adapt to new tasks Niklaus et al. (2024).
@@ 4.2 Frontier Models
We show results for frontier and reasoning models in Table 3. GPT-4o underperforms both of its peers Claude-Sonnet and Llama-3.1-405B. This is particularly unexpected, as models tend to favor their own completions Panickssery et al. (2024), and GEMBA-MQM is operated by GPT-4o. Claude-Sonnet demonstrates strong performance, competing closely with o1, the top-performing model. Surprisingly, o1-mini performs only on par with the other models at the smaller scale and even underperforms Claude-Haiku. Overall, Anthropic’s models are really strong, and even more so from a cost-to-performance perspective compared to reasoning models like o1.
@@ 4.6 Comparison Across Languages
In Figure 4, we compare the best models per category across language directions on CH-Law-Trans. Performance to and from German, French, and Italian is homogeneous across models. When translating from English to the other languages, all models perform worse than from the three main Swiss languages. Since the English source texts are already translations and are not legally binding, the federal translators may have applied less rigor in generating them, potentially resulting in lower quality and slight deviations. Anecdotally, the lawyers co-authoring this work confirm that the English source texts are occasionally less precise. So, the lower scores may also indicate that the judge model bases its grading on imperfect source text. Romansh is a low-resource language and only spoken by less than 50K people in Switzerland.777https://www.bfs.admin.ch/bfs/de/home/statistiken/bevoelkerung/sprachen-religionen/sprachen.html It is not supported by most translation models such as MADLAD-400. Surprisingly, o1 and Sonnet still perform very well when translating from Romansh to other languages. When translating to Romansh, all models’ quality drops off, sometimes sharply. Perhaps similar to humans, also for LLMs speaking or writing a language seems harder than understanding it.
@@ 5 Expert Evaluation
The experts were asked to assign a score between 0 and 100 to each translation. For this purpose, the experts were given a source text, its “gold translation” (official translation of the Swiss authorities) as a reference and a predicted translation. The scores only reflected the completeness and accuracy of the predicted translation, with less emphasis on readability and other stylistic attributes. To ensure consistency, the experts agreed on a point deduction system in advance and discussed certain borderline cases (annotation guidelines are in Appendix G). In total, 200 translations were assigned a score by at least two experts. Each expert assigned scores independently, without consulting the other annotators. For the expert agreement with judge metrics (see section 6) and for the evaluation of the best models (see subsection 5.2), we averaged the scores of the two annotators.
@@ 5.1 Inter-Annotator Agreement
The average Krippendorff’s was 0.56 for laws and 0.41 for headnotes. Agreement was generally higher for laws than for headnotes, likely because laws in our dataset were typically shorter. The moderate inter-annotator agreement suggests that, despite clear instructions, a certain degree of subjectivity was inherent in the task. In addition, we observed smaller differences between the individual language pairs, suggesting that not all annotators were perfectly aligned. However, disagreements tended to be minor and were rarely fundamental. In Figure 5 we show the absolute point difference between the two annotators evaluating the same samples. In almost half of the cases the two annotators completely agree and in 92% the difference is smaller than 30 points.
@@ 6.1 Setup
We use GPT-4o, GPT-4o-mini, Gemini-1.5-pro, and Gemini-1.5-flash in our judge model ablation. We also tested Claude Sonnet and Haiku as judges, but they failed to follow grading instructions.888They would insist on generating JSON output while we very clearly just asked for plain-text. The o1 and o1-mini models showed very low or even negative correlations with human judgments and are thus excluded. We randomly selected one few-shot example from the dev sets of laws, headnotes, and press releases. To ensure judge models saw diverse translation qualities, we chose models of varying strengths (Claude 3.5 Sonnet for laws, Mixtral-8x7B-Instruct-v0.1 for headnotes, and Qwen2.5-1.5B-Instruct for press releases). We used a simple prompt “Translate to target-language” to generate translations. Sample judgments per few-shot example were written by one lawyer author and double-checked by another. We tested two few-shot styles single (all examples in one language direction: fr-de) and diverse (law article en-it, headnote de-fr and press release fr-de). We ablated two user prompts with absolute grading (basic and detailed) and one with deduction grading similar to the codebook given to the human expert annotators (codebook). Judge prompts are in Appendix F.
@@
A: On average, the language directions of the few shot examples do not matter, but the best configuration uses diverse language directions. Across all 24 investigated configurations, there is no significant difference between the single and diverse language direction setup ( vs , ). However, the best configuration overall, uses diverse language directions.
@@
A: Our SwiLTra-Judge exhibits the highest correlation with human judgments among tested translation metrics. Figure 6 shows Spearman correlation with human judgments for sample-level metrics (this excludes BLEU and ChrF). As expected, METEOR and BERTScore perform poorly, with correlations below 0.2. Surprisingly, the recent GEMBA-MQM metric both underperforms BLEURT and XCOMET. Our SwiLTra-Judge is significantly better than the second-best metric XCOMET ( vs , ).
@@ 8 Conclusions
Our fine-tuned models are much stronger than the initial instruction-tuned open models they are based on, but they still under-perform large closed models. Future work could investigate techniques such as model merging Yang et al. (2024) to further improve and bring them closer to the frontier models. While we evaluated a large variety of models, we could not evaluate them all. Future work could investigate other promising models such as Grok.999https://x.ai/blog/grok-2 We took great care to validate our results with human expert studies. However, our resources were limited and we could not investigate certain languages (e.g., Romansh) and our sample sizes were still rather small. Future work could perform a more broad and in-depth human evaluation.
@@ Appendix E Judge Correlations
| gemini-1-5-flash-codebook-diverse-deduction | 0.504 [0.419, 0.587] | 0.504 ± 0.074 | 15.215 ± 2.216 | 11.100 ± 0.670 |
| XCOMET-XXL | 0.486 [0.403, 0.560] | 0.477 ± 0.093 | 14.877 ± 1.372 | 10.204 ± 0.748 |
| gpt-4o-mini-codebook-single-deduction | 0.470 [0.382, 0.549] | 0.477 ± 0.095 | 22.168 ± 3.064 | 16.944 ± 1.691 |
| gpt-4o-mini-codebook-diverse-deduction | 0.461 [0.374, 0.539] | 0.459 ± 0.094 | 22.527 ± 2.678 | 17.138 ± 1.113 |
| gemini-1-5-flash-codebook-single-deduction | 0.459 [0.373, 0.547] | 0.466 ± 0.069 | 16.049 ± 2.212 | 10.990 ± 0.720 |
| gpt-4o-codebook-single-deduction | 0.443 [0.357, 0.534] | 0.447 ± 0.070 | 29.020 ± 3.982 | 19.606 ± 1.951 |
| gpt-4o-codebook-diverse-deduction | 0.424 [0.335, 0.513] | 0.412 ± 0.044 | 30.902 ± 1.843 | 21.537 ± 0.555 |
| gpt-4o-detailed-single-absolute | 0.422 [0.337, 0.510] | 0.427 ± 0.052 | 35.940 ± 4.207 | 24.286 ± 3.382 |
| gpt-4o-mini-basic-single-absolute | 0.415 [0.333, 0.499] | 0.422 ± 0.067 | 26.736 ± 1.817 | 21.370 ± 2.036 |
| gpt-4o-basic-single-absolute | 0.411 [0.315, 0.499] | 0.411 ± 0.131 | 20.655 ± 2.947 | 14.596 ± 1.976 |
| gpt-4o-mini-detailed-single-absolute | 0.379 [0.287, 0.466] | 0.384 ± 0.069 | 35.437 ± 1.866 | 30.302 ± 1.718 |
| gpt-4o-basic-diverse-absolute | 0.379 [0.282, 0.468] | 0.383 ± 0.087 | 21.780 ± 2.249 | 13.550 ± 1.382 |
| gpt-4o-detailed-diverse-absolute | 0.377 [0.285, 0.466] | 0.380 ± 0.090 | 34.786 ± 3.711 | 22.995 ± 3.173 |
| bleurt_large | 0.356 [0.259, 0.446] | 0.364 ± 0.147 | 63.110 ± 5.225 | 58.102 ± 5.064 |
| gpt-4o-mini-detailed-diverse-absolute | 0.356 [0.259, 0.445] | 0.364 ± 0.097 | 36.480 ± 2.365 | 30.828 ± 2.387 |
| gpt-4o-mini-basic-diverse-absolute | 0.352 [0.262, 0.439] | 0.361 ± 0.048 | 28.363 ± 1.315 | 22.393 ± 1.347 |
| gemini-1-5-pro-basic-single-absolute | 0.305 [0.204, 0.400] | 0.295 ± 0.083 | 36.203 ± 3.618 | 22.993 ± 3.010 |
| gemini-1-5-pro-codebook-diverse-deduction | 0.302 [0.211, 0.404] | 0.298 ± 0.095 | 34.528 ± 3.882 | 20.580 ± 2.477 |
| gemini-1-5-pro-codebook-single-deduction | 0.293 [0.181, 0.391] | 0.292 ± 0.074 | 36.718 ± 2.663 | 21.816 ± 2.158 |
| GEMBA-MQM_gpt-4o | 0.290 [0.187, 0.383] | 0.289 ± 0.093 | 18.331 ± 1.743 | 12.698 ± 0.787 |
| gemini-1-5-flash-detailed-single-absolute | 0.279 [0.178, 0.365] | 0.275 ± 0.049 | 29.067 ± 3.070 | 18.709 ± 1.651 |
| gemini-1-5-flash-basic-single-absolute | 0.271 [0.171, 0.365] | 0.279 ± 0.110 | 33.283 ± 5.239 | 20.412 ± 3.287 |
| gemini-1-5-flash-basic-diverse-absolute | 0.258 [0.156, 0.351] | 0.249 ± 0.069 | 27.649 ± 5.852 | 16.756 ± 3.233 |
| gemini-1-5-pro-basic-diverse-absolute | 0.252 [0.154, 0.347] | 0.250 ± 0.082 | 36.544 ± 3.470 | 22.990 ± 2.684 |
| gemini-1-5-pro-detailed-diverse-absolute | 0.246 [0.142, 0.344] | 0.250 ± 0.097 | 38.316 ± 2.063 | 26.098 ± 1.148 |
| gemini-1-5-pro-detailed-single-absolute | 0.237 [0.135, 0.333] | 0.244 ± 0.079 | 38.618 ± 2.362 | 27.798 ± 1.977 |
| gemini-1-5-flash-detailed-diverse-absolute | 0.233 [0.131, 0.326] | 0.225 ± 0.055 | 30.091 ± 4.891 | 19.671 ± 3.346 |
| BERTScore-F | 0.164 [0.068, 0.260] | 0.170 ± 0.053 | 36.723 ± 1.340 | 31.523 ± 1.772 |
| meteor | 0.161 [0.063, 0.257] | 0.164 ± 0.125 | 34.170 ± 3.444 | 29.270 ± 3.191 |

### v2

@@ Abstract
In Switzerland legal translation is uniquely important due to the country’s four official languages and requirements for multilingual legal documentation. However, this process traditionally relies on professionals who must be both legal experts and skilled translators—creating bottlenecks and impacting effective access to justice. To address this challenge, we introduce SwiLTra-Bench, a comprehensive multilingual benchmark of over 180K aligned Swiss legal translation pairs comprising laws, headnotes, and press releases across all Swiss languages along with English, designed to evaluate LLM-based translation systems. Our systematic evaluation reveals that frontier models achieve superior translation performance across all document types, while specialized translation systems excel specifically in laws but under-perform in headnotes. Through rigorous testing and human expert validation, we demonstrate that while fine-tuning open SLMs significantly improves their translation quality, they still lag behind the best zero-shot prompted frontier models such as Claude-3.5-Sonnet. Additionally, we present SwiLTra-Judge, a specialized LLM evaluation system that aligns best with human expert assessments.
@@
Joel Niklaus Jakob Merane Luka Nenadic Sina Ahmadi Yingqiang Gao Cyrill A. H. Chevalley Claude Humbel Christophe Gösken Lorenzo Tanzi Thomas Lüthi Stefan Palombo Spencer Poff Boling Yang Nan Wu Matthew Guillod Robin Mamié Daniel Brunner Julio Pereyra Niko Grupen Harvey ETH Zurich Swiss Federal Supreme Court University of Zurich University of Basel University of Geneva University of Lausanne Canton of Solothurn Max Planck Institute for Research on Collective Goods Correspondence: joel@niklaus.ai
@@ 2.2 Dataset Splits
We first segment each dataset by a unique identifier (entire laws, entire headnotes, and entire press releases) to ensure that no single law, headnote, or press release is split across training, validation, and test. For laws, we prioritize examples for the validation and test splits that (1) have more language versions (to guarantee good multilingual coverage), (2) have an official abbreviation (since abbreviations are only set for those laws that are presumed to be cited frequently111https://www.bk.admin.ch/apps/gtr/de/index.html, which we consider a good proxy for practical importance, (3) have shorter text lengths to make evaluation faster and cheaper, and (4) have newer applicability dates so that more recent and multilingual laws are prioritized for validation and testing, resulting in a more realistic evaluation setting. For headnotes, we similarly prioritize those with more recent publication years for validation and test. Finally, for press releases, we focus on maximizing multilingual coverage by ensuring all validation and test examples are available in all present languages (German, French and Italian). The training sets contain all examples not held out for validation or testing.
@@ 2.3 Data Statistics
Table 1 presents the overall statistics of the three datasets included in SwiLTRa-Bench. We visualize the training set text lengths for the shortest levels used for training and evaluation in Figure B.1. For completeness, we show histograms for all levels in Figure B.2 and Figure B.3. To calculate these statistics, we used an NLTK222https://www.nltk.org word tokenizer, splitting sentences based on whitespace and punctuation.
@@ 3.1 Evaluation
To paint a representative picture of translation capabilities, we evaluate models across five main classes: 1) translation models, i.e., models specifically trained for translation tasks, 2) frontier models, i.e., large foundation models pre-trained on web-scale data and post-trained on diverse tasks, 3) reasoning models, i.e., models using significant resources at test time to improve output quality, 4) open models, i.e., typically small language models (SLMs) with publicly available weights, and 5) fine-tuned models, i.e., models specifically fine-tuned on SwiLTRa-Bench. We conducted our evaluation using the lighteval framework due to its ease of use and good support for custom metrics.333https://github.com/huggingface/lighteval
@@ 3.2 Fine-tuning
To provide an overview of the current open SLM landscape, we fine-tuned Gemma-2 2B and 9B Team et al. (2024), Llama 1B, 3B and 8B Grattafiori et al. (2024), Phi-3.5 mini and Phi-3 medium Abdin et al. (2024), and Qwen2.5 0.5B, 1.5B, 3B, 7B, 14B and 32B Team (2024) models on our dataset. We fine-tuned using Hugging Face transformers444https://github.com/huggingface/transformers and unsloth555https://github.com/unslothai/unsloth using 4-bit quantization and 8bit AdamW Loshchilov and Hutter (2019); Dettmers et al. (2022) on a single 80GB NVIDIA H100 GPU. We used rank stabilized LoRA Hu et al. (2021); Kalajdzievski (2023) with rank 16 and alpha 16. We trained with the model’s native chat template on sequence length 512, covering more than 99% of the training dataset and truncating the rest. The instruction template was simply:
{source language}: {source text} {target language}: {target text}
We trained on the entire training set for maximal coverage of the data. We used packing, weight decay 0.01, batch size 128 and early stopping with patience 3. In most cases, the lowest evaluation loss is reached after exactly 1 epoch. We used a linear learning rate schedule with 1000 warmup steps and learning rate . We manually tuned the learning rate ( - ), weight decay (0.01, 0.1), label smoothing (factor 0, 0.01, 0.1) and LoRA rank (16, 128). We used the train and validation sets of the Law and Headnote translations on the lowest (shortest) levels, i.e., the paragraph and text levels. For all fine-tuned models, we used the instruction-tuned variant since they have shown to better adapt to new tasks Niklaus et al. (2024).
@@ 4 Results and Analysis
4.1 Statistical Significance Testing
The model comparisons presented throughout this section are descriptive in nature. When we state that one model “outperforms” or “outcompetes” another, we are describing observed differences in the reported metrics; we do not claim statistical significance unless explicitly stated. We only apply null hypothesis significance testing (NHST) in those instances where we explicitly report statistically significant results. In such cases, we provide information about the tests used, statistics obtained, and effect sizes.
@@ 4.3 Frontier Models
We show results for frontier and reasoning models in Table 3. GPT-4o underperforms both of its peers Claude-3.5-Sonnet and Llama-3.1-405B. This is particularly unexpected, as models tend to favor their own completions Panickssery et al. (2024), and GEMBA-MQM is operated by GPT-4o. Claude-3.5-Sonnet demonstrates strong performance, competing closely with o1, the top-performing model. Surprisingly, o1-mini performs only on par with the other models at the smaller scale and even underperforms Claude-3.5-Haiku. Overall, Anthropic’s models are really strong, and even more so from a cost-to-performance perspective compared to reasoning models like o1.
@@ 4.7 Comparison Across Languages
In Figure 4, we compare the best models per category across language directions on CH-Law-Trans. Performance to and from German, French, and Italian is homogeneous across models. When translating from English to the other languages, all models perform worse than from the three main Swiss languages. Since the English source texts are already translations and are not legally binding, the federal translators may have applied less rigor in generating them, potentially resulting in lower quality and slight deviations. Anecdotally, the lawyers co-authoring this work confirm that the English source texts are occasionally less precise. So, the lower scores may also indicate that the judge model bases its grading on imperfect source text.
Romansh is a low-resource language and only spoken by less than 50K people in Switzerland.666https://www.bfs.admin.ch/bfs/de/home/statistiken/bevoelkerung/sprachen-religionen/sprachen.html As our dataset consists of the entire data readily available for federal laws, Supreme Court headnotes, and Supreme Court press releases in Switzerland, it is not possible to extend the Romansh coverage. Adding additional data sources is not a viable option due to likely lower data quality and limited sentence level alignment across languages. Romansh is not supported by most translation models such as MADLAD-400. In our opinion, the fact that these models don’t support Romansh at the moment highlights the value of our dataset for low-resource languages. Indeed, our dataset now enables teams building translation models to train and evaluate on >160K and >8K Romansh translation pairs respectively. Surprisingly, o1 and Sonnet still perform very well when translating from Romansh to other languages. When translating to Romansh, all models’ quality drops off, sometimes sharply. Perhaps similarly to humans, also for LLMs speaking or writing a language seems harder than understanding it.
@@ 5 Expert Evaluation
The experts were asked to assign a score between 0 and 100 to each translation. For this purpose, the experts were given a source text, its “gold translation” (official translation of the Swiss authorities) as a reference and a predicted translation. The scores only reflected the completeness and accuracy of the predicted translation, with less emphasis on readability and other stylistic attributes. To ensure consistency, the experts agreed on a point deduction system in advance and discussed certain borderline cases (annotation guidelines are in Appendix G). In total, 200 translations were assigned a score by exactly two experts. Each expert assigned scores independently, without consulting the other annotators. For the expert agreement with judge metrics (see section 6) and for the evaluation of the best models (see subsection 5.2), we averaged the scores of the two annotators.
@@ 5.1 Inter-Annotator Agreement
The average Krippendorff’s was 0.56 for laws and 0.41 for headnotes. Agreement was generally higher for laws than for headnotes across most language pairs and tasks. This consistent pattern is unlikely due to a calibration issue and instead suggests that the headnote task inherently allows for more subjectivity. This is likely due to headnotes being longer on average and containing greater interpretive nuance, which can lead annotators to miss or accord different importance to different details. We also hypothesize that laws inherently exhibit a more structured and clearer pattern, making them easier to translate and evaluate.
The moderate inter-annotator agreement suggests that, despite clear instructions, a certain degree of subjectivity was inherent in the task. In addition, we observed smaller differences between the individual language pairs, suggesting that not all annotators were perfectly aligned. However, disagreements tended to be minor and were rarely fundamental. In Figure 5 we show the absolute point difference between the two annotators evaluating the same samples. In almost half of the cases the two annotators completely agree and in 92% the difference is smaller than 30 points.
@@ 6.1 Setup
We use GPT-4o, GPT-4o-mini, Gemini-1.5-pro, and Gemini-1.5-flash in our judge model ablation. We also tested Claude Sonnet and Haiku as judges, but they failed to follow grading instructions.777They would insist on generating JSON output while we very clearly just asked for plain-text. The o1 and o1-mini models showed very low or even negative correlations with human judgments and are thus excluded. We randomly selected one few-shot example from the dev sets of laws, headnotes, and press releases. To ensure judge models saw diverse translation qualities, we chose models of varying strengths (Claude 3.5 Sonnet for laws, Mixtral-8x7B-Instruct-v0.1 for headnotes, and Qwen2.5-1.5B-Instruct for press releases). We used a simple prompt “Translate to target-language” to generate translations. Sample judgments per few-shot example were written by one lawyer author and double-checked by another. We tested two few-shot styles single (all examples in one language direction: fr-de) and diverse (law article en-it, headnote de-fr and press release fr-de). We ablated two user prompts with absolute grading (basic and detailed) and one with deduction grading similar to the codebook given to the human expert annotators (codebook). Judge prompts are in Appendix F.
@@
A: On average, the language directions of the few shot examples do not matter, but the best configuration uses diverse language directions. Across all 24 investigated configurations, there is only minimal difference between the single and diverse language direction setup ( vs ). However, the best configuration overall, uses diverse language directions.
@@
A: Our SwiLTra-Judge exhibits the highest correlation with human judgments among tested translation metrics. Figure 6 shows Spearman correlation with human judgments for sample-level metrics (this excludes BLEU and ChrF). As expected, METEOR and BERTScore perform poorly, with correlations below 0.2. Surprisingly, the recent GEMBA-MQM metric both underperforms BLEURT and XCOMET. Our SwiLTra-Judge-Single (gemini-1-5-flash-codebook-diverse-deduction) is significantly better than the second-best metric XCOMET ( vs , ). Our SwiLTra-Judge-Ensemble (a simple mean score from GPT-4o-mini and Gemini 1.5 Flash with codebook style and both single and diverse few-shot setup) is again significantly better than our SwiLTra-Judge-Single ( vs , ). To compute both p-values, we have performed an unpaired t-test with each.
@@ 8 Conclusions
Our fine-tuned models are much stronger than the initial instruction-tuned open models they are based on, but they still under-perform large closed models. Future work could investigate techniques such as model merging Yang et al. (2024) to further improve and bring them closer to the frontier models. While we evaluated a large variety of models, we could not evaluate them all. Future work could investigate other promising models such as Grok.888https://x.ai/blog/grok-2 We took great care to validate our results with human expert studies. However, our resources were limited and we could not investigate certain languages (e.g., Romansh) and our sample sizes were still rather small. Future work could perform a more broad and in-depth human evaluation.
@@
Acknowledgments
We thank the anonymous reviewers for their constructive feedback. Luka Nenadic’s research is funded by SNSF grant No. 10002634.
@@ Appendix E Judge Correlations
| judge-ensemble | 0.536 [0.453, 0.608] | 0.533 ± 0.080 | 16.090 ± 1.424 | 12.979 ± 0.882 |
| gemini-1-5-flash-codebook-diverse-deduction | 0.506 [0.424, 0.586] | 0.504 ± 0.074 | 15.215 ± 2.216 | 11.100 ± 0.670 |
| XCOMET-XXL | 0.484 [0.403, 0.567] | 0.477 ± 0.093 | 14.877 ± 1.372 | 10.204 ± 0.748 |
| gpt-4o-mini-codebook-single-deduction | 0.474 [0.387, 0.551] | 0.477 ± 0.095 | 22.168 ± 3.064 | 16.944 ± 1.691 |
| gemini-1-5-flash-codebook-single-deduction | 0.461 [0.374, 0.542] | 0.466 ± 0.069 | 16.049 ± 2.212 | 10.990 ± 0.720 |
| gpt-4o-mini-codebook-diverse-deduction | 0.459 [0.378, 0.538] | 0.459 ± 0.094 | 22.527 ± 2.678 | 17.138 ± 1.113 |
| gpt-4o-codebook-single-deduction | 0.444 [0.352, 0.538] | 0.447 ± 0.070 | 29.020 ± 3.982 | 19.606 ± 1.951 |
| gpt-4o-codebook-diverse-deduction | 0.427 [0.334, 0.516] | 0.412 ± 0.044 | 30.902 ± 1.843 | 21.537 ± 0.555 |
| gpt-4o-detailed-single-absolute | 0.418 [0.330, 0.501] | 0.427 ± 0.052 | 35.940 ± 4.207 | 24.286 ± 3.382 |
| gpt-4o-mini-basic-single-absolute | 0.413 [0.320, 0.497] | 0.422 ± 0.067 | 26.736 ± 1.817 | 21.370 ± 2.036 |
| gpt-4o-basic-single-absolute | 0.411 [0.320, 0.499] | 0.411 ± 0.131 | 20.655 ± 2.947 | 14.596 ± 1.976 |
| gpt-4o-detailed-diverse-absolute | 0.378 [0.284, 0.464] | 0.380 ± 0.090 | 34.786 ± 3.711 | 22.995 ± 3.173 |
| gpt-4o-mini-detailed-single-absolute | 0.378 [0.282, 0.471] | 0.384 ± 0.069 | 35.437 ± 1.866 | 30.302 ± 1.718 |
| gpt-4o-basic-diverse-absolute | 0.376 [0.281, 0.467] | 0.383 ± 0.087 | 21.780 ± 2.249 | 13.550 ± 1.382 |
| gpt-4o-mini-detailed-diverse-absolute | 0.358 [0.259, 0.450] | 0.364 ± 0.097 | 36.480 ± 2.365 | 30.828 ± 2.387 |
| bleurt_large | 0.356 [0.261, 0.445] | 0.364 ± 0.147 | 63.110 ± 5.225 | 58.102 ± 5.064 |
| gpt-4o-mini-basic-diverse-absolute | 0.354 [0.263, 0.447] | 0.361 ± 0.048 | 28.363 ± 1.315 | 22.393 ± 1.347 |
| gemini-1-5-pro-basic-single-absolute | 0.306 [0.209, 0.404] | 0.295 ± 0.083 | 36.203 ± 3.618 | 22.993 ± 3.010 |
| gemini-1-5-pro-codebook-diverse-deduction | 0.303 [0.200, 0.397] | 0.298 ± 0.095 | 34.528 ± 3.882 | 20.580 ± 2.477 |
| GEMBA-MQM_gpt-4o | 0.293 [0.189, 0.383] | 0.289 ± 0.093 | 18.331 ± 1.743 | 12.698 ± 0.787 |
| gemini-1-5-pro-codebook-single-deduction | 0.292 [0.195, 0.392] | 0.292 ± 0.074 | 36.718 ± 2.663 | 21.816 ± 2.158 |
| gemini-1-5-flash-detailed-single-absolute | 0.281 [0.190, 0.375] | 0.275 ± 0.049 | 29.067 ± 3.070 | 18.709 ± 1.651 |
| gemini-1-5-flash-basic-single-absolute | 0.270 [0.179, 0.359] | 0.279 ± 0.110 | 33.283 ± 5.239 | 20.412 ± 3.287 |
| gemini-1-5-flash-basic-diverse-absolute | 0.255 [0.157, 0.351] | 0.249 ± 0.069 | 27.649 ± 5.852 | 16.756 ± 3.233 |
| gemini-1-5-pro-basic-diverse-absolute | 0.252 [0.153, 0.351] | 0.250 ± 0.082 | 36.544 ± 3.470 | 22.990 ± 2.684 |
| gemini-1-5-pro-detailed-diverse-absolute | 0.247 [0.147, 0.345] | 0.250 ± 0.097 | 38.316 ± 2.063 | 26.098 ± 1.148 |
| gemini-1-5-pro-detailed-single-absolute | 0.235 [0.137, 0.334] | 0.244 ± 0.079 | 38.618 ± 2.362 | 27.798 ± 1.977 |
| gemini-1-5-flash-detailed-diverse-absolute | 0.231 [0.130, 0.328] | 0.225 ± 0.055 | 30.091 ± 4.891 | 19.671 ± 3.346 |
| BERTScore-F | 0.163 [0.066, 0.268] | 0.170 ± 0.053 | 36.723 ± 1.340 | 31.523 ± 1.772 |
| meteor | 0.160 [0.057, 0.259] | 0.164 ± 0.125 | 34.170 ± 3.444 | 29.270 ± 3.191 |